# The SNEWS 2.0 Alert Software for the Coincident Detection of Neutrinos from Core-Collapse Supernovae

### v1

@@ 1 SNEWS Introduction
The Supernova Early Warning System (SNEWS) is a global network of detectors that aims to provide early warning of nearby core-collapse supernova (CCSN) explosions. The neutrino signal from a core-collapse supernova, which significantly increases at the point of explosion, provides a valuable early warning for such celestial events [1]. [^1] Neutrinos are emitted in vast quantities from a star as it collapses, reaching a peak coinciding with the supernova explosion. This signal precedes the optical visibility of the supernova, making it crucial for early detection. SNEWS aims to quickly distribute the early warnings across a network of neutrino and dark-matter detectors. Each detector is independently sensitive to CCSN neutrinos. Taken together, the combined signal provides richer information than any single experiment could achieve alone.
@@
This network continues to operate and will continue to operate until the new SNEWS 2.0 is well established. However, there are several reasons to upgrade. First, the original code, written 25 years ago for now obsolete VMS, IRIX, and AIX computers, is hard to port to new experiments. In addition to maintainability, the current proliferation of MMA projects has changed the community’s perception of alert publishing. When SNEWS started, one false alarm would have been the death of the project. Now, people realize that to maximize results, false alarms can be tolerated, provided that confidence levels are provided so consumers can choose their own appropriate action. Furthermore, with the increased size of modern experiments, the once statistically intractable problem of using neutrino arrival times to point back to an area on the sky is now possible, with a similar uncertainty to the LIGO/VIRGO gravitational wave skymaps [5, 6, 7] that enabled the identification of electromagnetic counterparts. Finally, a wide range of automated follow-up observatories are now active and chasing after gamma-ray bursts, gravitational waves, and high-energy neutrinos. Thus, the SNEWS 2.0 effort [3] to address these changes.
@@ 2.3 Subscribing
snews_pt offers a user-friendly and efficient method for subscribing to the SNEWS alert channel. Notably, this subscription process is accessible to a wide user base, extending beyond member experiments to include any individuals or observatory with Kafka credentials. These users can subscribe and actively monitor the incoming alerts using the Subscriber class, accessible through the API or the terminal.
@@ 2.5 Retraction and Skipped Heartbeat Warnings
Heartbeats play a crucial role in the overall functionality of the early warning system. Through the heartbeats, the server running the SNEWS Coincidence System can actively monitor and alert users in case of a lost connection at any given point. While the server anticipates receiving heartbeats from all detectors, it does not impose a specific frequency requirement. Once the server accumulates enough heartbeats, it calculates the heartbeat frequency and assigns a value of $\mu\pm\sigma$ to the corresponding experiment. With the continuous accumulation of heartbeats, statistical accuracy improves, and the expected frequency is dynamically updated. With the established frequency, the server anticipates receiving the next heartbeat within this time interval. The server conducts this verification process every minute, and if no new heartbeat is registered after the $\mu+3\sigma$ time has elapsed, it generates a warning email. This email includes the collected statistics and is sent to the pre-registered responsible email addresses associated with that specific experiment. Upon the absence of new heartbeats and the generation of a warning email, the checks for this particular experiment are temporarily halted. The system resumes the frequency computation once new heartbeats start accumulating, ensuring that the calculations are based on the most recent data. This adaptive approach allows the system to dynamically adjust to changing conditions and provides accurate frequency assessments for effective monitoring and alert handling.
@@ 4 Firedrills
SNEWS holds periodic distributed tests called “firedrills” to test that both snews_pt and snews_cs are operating as intended, to validate the pointing capabilities, and to test system updates.
@@
For the development team, firedrills allow a wider audience to give feedback on the snews_pt and the alert protocols of the snews_cs. Furthermore, edge cases, such as staggered coincidences or updating an alert when the detector retracts a message, are tested during firedrills. Frequent testing helps the participating members of SNEWS to familiarize themselves with deploying snews_pt. A typical firedrill involves SNEWS members participating in a virtual call to test various alert scenarios. These scenarios include a simple coincidence, where all detectors register the supernova flux simultaneously, a staggered coincidence involving a subset of detectors with delayed data transmission, and a coincidence with retractions and detector updates. Firedrills provide a crucial learning experience for the development team to enhance the software and refine the implementation of these systems within SNEWS. Notably, the most significant updates to both snews_pt and snews_cs were implemented after thorough feedback by the SNEWS community following a firedrill. For instance, during a firedrill in October 2022, participating SNEWS members contributed the following improvements: changed the snews_pt logic to exclude testing messages for detectors; established a proper naming convention for detectors; enabled messages to include time uncertainty of the neutrino signal; and updated the alert statistics following an update or retraction.
@@ 5 Conclusion
By leveraging a more flexible and versatile code base, SNEWS member experiments can more easily publish and receive alerts to potential galactic supernovae. SNEWS 2.0 has developed a Python-based framework based on SCIMMA’s HOPSKOTCH system for communications with neutrino detectors and the world. This Python framework consists of two parts: snews_pt allows experiments to communicate their potential supernova alerts, and snews_cs receives those alerts, forming coincidence alerts if two or more unique signals are within a $10$-second window. Once coincident signals are detected, a prompt alert is published to the MMA community and the public. This framework implements a wide range of plugins to receive the alert and distribute it in various required formats, including email, GCN alerts, or directly integrating with snews_pt to feed custom code back into an experiment or observatory. Each software is designed to be sustainable and scalable, allowing SNEWS 2.0 to incorporate additional functionalities such as pointing or source significance stacking. With SNEWS 2.0, the MMA community will receive to receive an alert prompt enough to adapt the orientation of telescopes and antennas accordingly, assuring the most sensitive detection of the burst, providing a unique chance to learn about currently un-observed details of the explosion process. Additionally, the capability to reduce the uncertainty on the start of the burst can significantly aid the chances to detect gravitational waves from a CCSN [14].
@@ Acknowledgement
This work is supported by the National Science Foundation “Windows on the Universe: The Era of Multi-Messenger Astrophysics” program: “WoU-MMA: Collaborative Research: Advancing the SuperNova Early Warning System" through grants 2209449, 2209451, and 2209534. We also acknowledge the support from the Science and Technology Facilities Council (STFC), United Kingdom.

### v2

@@ 1 SNEWS Introduction
The Supernova Early Warning System (SNEWS) is a global network of detectors that aims to provide early warning of nearby core-collapse supernova (CCSN) explosions. The neutrino signal from a core-collapse supernova, which significantly increases at the point of explosion, provides a valuable early warning for such celestial events [1]. [^1] Neutrinos are emitted in vast quantities from a star as it collapses, reaching a peak coinciding with the supernova explosion. This signal precedes the optical visibility of the supernova, making it crucial for early detection. SNEWS aims to quickly distribute the early warnings across a network of neutrino and dark matter detectors. Each detector is independently sensitive to CCSN neutrinos. Taken together, the combined signal provides richer information than any single experiment could achieve alone.
@@
This network continues to operate and will continue to operate until the new SNEWS 2.0 is well established. However, there are several reasons to upgrade. First, the original code, written 25 years ago for now obsolete VMS, IRIX, and AIX computers, is hard to port to new experiments. In addition to maintainability, the current proliferation of MMA projects has changed the community’s perception of alert publishing. When SNEWS started, one false alarm would have been the death of the project. Now, people realize that to maximize results, false alarms can be tolerated, provided that confidence levels are provided so consumers can choose their own appropriate action. Furthermore, with the increased size of modern experiments, the once statistically intractable problem of using neutrino arrival times to point back to an area in the sky is now possible, with a similar uncertainty to the LIGO/VIRGO gravitational wave skymaps [5, 6, 7] that enabled the identification of electromagnetic counterparts. Finally, a wide range of automated follow-up observatories are now active and chasing after gamma-ray bursts, gravitational waves, and high-energy neutrinos. Thus, the SNEWS 2.0 effort [3] to address these changes.
@@ 2.3 Subscribing
snews_pt offers a user-friendly and efficient method for subscribing to the SNEWS alert channel. Notably, this subscription process is accessible to a wide user base, extending beyond member experiments to include any individuals or observatories with Kafka credentials. These users can subscribe and actively monitor the incoming alerts using the Subscriber class, accessible through the API or the terminal.
@@ 2.5 Retraction and Skipped Heartbeat Warnings
Heartbeats play a crucial role in the overall functionality of the early warning system. Through the heartbeats, the server running the SNEWS Coincidence System can actively monitor and alert users in case of a lost connection at any given point. While the server anticipates receiving heartbeats from all detectors, it does not impose a specific frequency requirement. Once the server accumulates enough heartbeats, it calculates the heartbeat frequency and assigns a value of $\mu\pm\sigma$ to the corresponding experiment. With the continuous accumulation of heartbeats, statistical accuracy improves, and the expected frequency is dynamically updated. With the established frequency, the server anticipates receiving the next heartbeat within this time interval. The server conducts this verification process every minute, and if no new heartbeat is registered after the $\mu+3\sigma$ time has elapsed, it generates a warning email. This email includes the collected statistics and is sent to the pre-registered responsible email addresses associated with that specific experiment. Upon the absence of new heartbeats and the generation of a warning email, the checks for this particular experiment are temporarily halted. The system resumes the frequency computation once new heartbeats start accumulating, ensuring that the calculations are based on the most recent data. This adaptive approach allows the system to adjust to changing conditions dynamically and provides accurate frequency assessments for effective monitoring and alert handling.
@@ 4 Firedrills
SNEWS conducts periodic distributed tests called “firedrills” to test that both snews_pt and snews_cs are operating as intended, to validate the pointing capabilities, and to test system updates.
@@
For the development team, firedrills allow a wider audience to give feedback on the snews_pt and the alert protocols of the snews_cs. Furthermore, edge cases, such as staggered coincidences or updating an alert when the detector retracts a message, are tested during firedrills. Frequent testing helps participating SNEWS members familiarize themselves with the deployment of snews_pt.
A typical firedrill involves SNEWS members participating in a virtual call to test various alert scenarios. These scenarios include a simple coincidence, where all detectors register the supernova flux simultaneously, a staggered coincidence involving a subset of detectors with delayed data transmission, and a coincidence with retractions and detector updates.
Firedrills provide a crucial learning experience for the development team to enhance the software and refine the implementation of these systems within SNEWS. Notably, the most significant updates to both snews_pt and snews_cs were implemented after thorough feedback by the SNEWS community following a firedrill. For instance, during a firedrill in October 2022, participating SNEWS members contributed the following improvements: changed the snews_pt logic to exclude testing messages for detectors; established a proper naming convention for detectors; enabled messages to include time uncertainty of the neutrino signal; and updated the alert statistics following an update or retraction.
@@ 5 Conclusion
By leveraging a more flexible and versatile code base, SNEWS member experiments can more easily publish and receive alerts to potential galactic supernovae. SNEWS 2.0 has developed a Python-based framework based on SCIMMA’s HOPSKOTCH system for communications with neutrino detectors and the world. This Python framework consists of two parts: snews_pt allows experiments to communicate their potential supernova alerts, and snews_cs receives those alerts, forming coincidence alerts if two or more unique signals are within a $10$-second window. Once coincident signals are detected, a prompt alert is published to the MMA community and the public. This framework implements a wide range of plugins to receive the alert and distribute it in various required formats, including email, GCN alerts, or directly integrating with snews_pt to feed custom code back into an experiment or observatory. Each software is designed to be sustainable and scalable, allowing SNEWS 2.0 to incorporate additional functionalities such as pointing or source significance stacking. With SNEWS 2.0, the MMA community will receive an alert prompt enough to adapt the orientation of telescopes and antennas accordingly, assuring the most sensitive detection of the burst, providing a unique chance to learn about currently un-observed details of the explosion process. Additionally, the capability to reduce the uncertainty on the start of the burst can significantly aid the chances to detect gravitational waves from a CCSN [14].
@@ Acknowledgement
This work is supported by the National Science Foundation “Windows on the Universe: The Era of Multi-Messenger Astrophysics” program: “WoU-MMA: Collaborative Research: Advancing the SuperNova Early Warning System" through grants 2209449, 2209451, and 2209534 as well as NSF AST-2206532. We also acknowledge the support from the Science and Technology Facilities Council (STFC), United Kingdom.